OpenStudy (anonymous):

Find the vertex of the graph of the function. f(x) = (x - 5)2 + 10

OpenStudy (anonymous):

Okay, whats the derivative?

OpenStudy (loser66):

ok, make the thing on the same way: the vertex form of a parabola is y = a (x - h )^2 + k your problem is y = 1 ( x - 5 )^2 + 10 " Si" the same form?

OpenStudy (anonymous):

-5 and 10 right?

OpenStudy (anonymous):

okay